8. Stacey wants to double her money in 5 years. What annual rate of return must she find in an investment to reach her goal? I/Y PV PMT FV
Answer #1

Given,

N = 5 years

Let Present Value = X

PMT = 0

FV = 2X

We need to calculate interest rate I/Y

Using Calculator, enter n = 5, PMT = 0, PV = -1, FV = 2,
Press CPT I/Y to get I/Y = 14.87%

Manual Calculation

FV = PV(1+r)n

=> 2X = X(1+r)5

=> r = 14.87%